OpenStudy (anonymous):

I'm so horrible at these, pleaseeeee help! One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single?

OpenStudy (anonymous):

x = 0.6x + 384 ^ x by itself is really the same as 1x 1x = 0.6x + 384 0.4x = 384 x = 960 Lets check our work: x = 0.6x + 384 (960) = 0.6(960) + 384 960 = 576 + 384 960 = 960 so we are correct :)

OpenStudy (anonymous):

@123456789987654321 do you get it?

OpenStudy (anonymous):

960 is the total downloads. question is asking to find Qsam downloads = 60 % of 960 = .6 * 960 = 576 :)

OpenStudy (anonymous):

yeaa i thinkk sooo thanksssss

OpenStudy (anonymous):

^ yeah i read it wrong ok so 960 is the total 960 * .6 = 576